# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v1

@@ 2.2. Relation to the periodic polymer
4. Zero-temperature analogues: proofs of Theorems 1.8 and 1.9
Lemma 4.1.
The limit as $\beta\to\infty$ of $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ is $G^{\mathbf{W}}(V\mid U)$, the (multi-path)-LPP time between $U$ and $V$. Further, this convergence is uniform as a function of $\mathbf{W}$.
Proof.
By Lemma A.1, this converges uniformly, since there are finitely many multi-paths between any two sets of points. ∎
Proof of Theorem 1.8..
$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (4.1)
$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (4.2)
By Lemma 4, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$. We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (4.3)
$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}$
We next show the braid relations pass to zero-temperature.
Proof of Theorem 1.9..
Defining $\mathcal{P}_{k}^{\beta}$ as in (4.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6. As noted in the proof of Theorem 1.8, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly. This immediately applies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
5. Burke properties and permutation invariance
In this section, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment and provide geometric and exponential analogues of other Burke property for the zero-temperature discrete periodic Pitman transform.
@@ 5. Burke properties and permutation invariance
The joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
@@
where we used Lemmas A.4-A.5 in the penultimate equality. Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This follows from Lemma A.7. ∎
Proposition 5.2 (Geometric Burke property).
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$
Proof.
Proposition 5.3 (Exponential Burke property).
Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences such that $X_{1,i}\sim\text{Exp}(a_{1}+b_{i})$ and $X_{2,i}\sim\text{Exp}(a_{2}+b_{i})$ for all $i\in\mathbb{Z}_{N}$ where $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$
Proof.
For $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$, define
$X_{m,i}^{\prime}\sim\text{Geom}\Biggl{(}\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)\Biggr{)}$
$\frac{1}{n}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{T}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)},\quad\text{and}\quad\frac{1}{n}\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{D}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)}$
By Proposition 5.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$. Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps. We then have that
5.1. Proofs of Theorems 1.10 and 1.11
We first prove a lemma that connects the condition stated in Theorem 1.10 to the condition in Theorem 1.4.
@@
for disjoint subsequences $k_{\ell_{1}},\ldots,k_{\ell_{r}}$ and $k_{p_{1}},\ldots,k_{p_{n}}$ with $r+n=m$. Since $\sigma_{1}$ only permutes elements of $\mathbb{Z}_{<x}$, we have that $k_{\ell_{t}}+1<x$ for $1\leq t\leq r$. Since $\sigma_{2}$ only permutes elements of $\mathbb{Z}_{\geq x}$, $k_{p_{t}}+1>x$ for $1\leq t\leq n$. Hence, $x\neq k_{\ell}+1$ for $1\leq\ell\leq m$. ∎
@@
For an arbitrary finite permutation $\sigma$, decompose it as a product of nearest-neighbor transpositions $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$. Take this decomposition to have the minimal number of terms. Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}$, as in Theorem 1.4. By Lemma 5.4, the condition (1.12) and the minimality of the decomposition ensures that the conditions of Theorem 1.4 are met, so
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)=Z^{\mathbf{W}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
Furthermore, by successively applying the Burke property in Proposition 5.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$. This completes the proof.
In the last-passage percolation setting, the proof is the same, instead using the invariance statement in Theorem 1.8 and the Burke properties in Propositions 5.2 and 5.3. ∎

### v2

@@ 2.2. Relation to the periodic polymer
4. Burke property and permutation invariance
In this section, we prove Theorems 1.7 and 1.8. We first need two intermediate lemmas. In the first, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment.
@@ 4. Burke property and permutation invariance
Using Lemmas A.4-A.5 in the penultimate equality below, the joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
@@
Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This follows from Lemma A.7. ∎
We next prove a lemma that connects the condition stated in Theorem 1.7 to the condition in Theorem 1.4.
@@
for disjoint subsequences $k_{\ell_{1}},\ldots,k_{\ell_{r}}$ and $k_{p_{1}},\ldots,k_{p_{n}}$ with $r+n=m$. Since $\sigma_{1}$ only permutes elements of $\mathbb{Z}_{<x}$, we have that $k_{\ell_{t}}+1<x$ for $1\leq t\leq r$. Since $\sigma_{2}$ only permutes elements of $\mathbb{Z}_{\geq x}$, we have $k_{p_{t}}+1>x$ for $1\leq t\leq n$. Hence, $x\neq k_{\ell}+1$ for $1\leq\ell\leq m$. ∎
@@
For an arbitrary finite permutation $\sigma$, decompose it as a product of nearest-neighbor transpositions $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$. Take this decomposition to have the minimal number of terms. Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$, as in Theorem 1.4. By Lemma 4.2, the condition (1.11) and the minimality of the decomposition ensures that the conditions of Theorem 1.4 are met, so
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U)=Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
Furthermore, by successively applying the Burke property in Proposition 4.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$. This completes the proof. ∎